# The Ginsburg–Sands theorem and computability theory

### v1

@@ 7. Separating $\mathsf{GST}_{1}$ from $\mathsf{RT}^{2}_{2}$
$\mathcal{P}_{e}$ ::: $\Phi^{P_{0}}_{e}$ is not the characteristic function of an infinite set or
not a non-hyperimmunity witness for $\overline{S}$.
We call requirements of the former kind the $\mathcal{L}$ requirements, and those of the latter kind the $\mathcal{P}$ requirements. It is clear that satisfying all $\mathcal{L}$ and $\mathcal{P}$ requirements ensures that $S$, $L_{0}$, and $P_{0}$ will have the desired properties. We assign a priority ordering to the requirements of type $\omega$, with $\mathcal{P}_{e}$ stronger than $\mathcal{L}_{e}$, and $\mathcal{L}_{e}$ stronger than $\mathcal{P}_{e+1}$, for all $e\in\omega$. Requirements will act at certain stages, never more than one requirement per stage, in a way that we will make precise below. When some $\mathcal{L}_{e}$ acts, it will impose a restraint $r_{e}$ that may be cancelled later and redefined later still; this will be used to control which elements can be enumerated into $S$ by lower priority $\mathcal{P}$ requirements. At each stage $s$, we also let $m_{s}$ be the maximum of all restraints (including cancelled or redefined ones) imposed at any point in the construction prior to stage $s$, as well as all numbers enumerated into $S$ so far. We say a requirement $\mathcal{L}_{e}$ requires attention at stage $s$ if there exist numbers $x_{1}>x_{0}\geq m_{s}$ such that $\Phi_{e}^{L_{0}}(x_{0})\operatorname{\downarrow}=\Phi_{e}^{L_{0}}(x_{1})\operatorname{\downarrow}=1$. Notice that since $L_{0}$ is low, determining if $\mathcal{L}_{e}$ requires attention is uniformly computable in $\emptyset^{\prime}$. For each $s$, let $T_{-1,s}$ be the set of all $\sigma\in T$ of length at most $s$. Now for each $e\leq s$, define a tree $T_{e,s}$ by induction. Let $U$ be the set of all $\sigma\in T_{e-1,s}$ such that there do not exist numbers $x_{0}<x_{1}\leq s$, larger than all active restraints at stage $s$ imposed by any $\mathcal{L}$ requirements stronger than $\mathcal{P}_{e}$, such that $\Phi_{e}^{\sigma}(x_{0})\operatorname{\downarrow}=\Phi_{e}^{\sigma}(x_{1})\operatorname{\downarrow}=1$. If $U$ contains a string of length $s$, we let $T_{e,s}=U$. Otherwise, we let $T_{e,s}=T_{e-1,s}$ and we say $\mathcal{P}_{e}$ requires attention at stage $s$. Since $T_{e,s}$ is uniformly computable in $\emptyset^{\prime}$, so is determining if $\mathcal{P}_{e}$ requires attention. This brings us to the construction itself. Initially, we let $S$ be empty and let no restraints be imposed. (In particular, $m_{0}=0$.) Now suppose we are at stage $s$. If it exists, fix the highest priority requirement of index $e\leq s$ that requires attention at stage $s$ and that has not acted since the last stage, if any, at which a higher priority $\mathcal{P}$ requirement acted. If this requirement is $\mathcal{L}_{e}$, we fix the least witnessing $x_{1}>x_{0}\geq m_{s}$ and impose the restraint $r_{e}=x_{1}$. We then say that $\mathcal{L}_{e}$ has acted. If, instead, this requirement is $\mathcal{P}_{e}$, we proceed as follows. By definition, for each $\sigma\in T_{e,s}$ of length $s$ there exist $x_{0}<x_{1}\leq s$, larger than all active restraints imposed by higher priority $\mathcal{L}$ requirements, such that $\Phi_{e}^{\sigma}(x_{0})\operatorname{\downarrow}=\Phi_{e}^{\sigma}(x_{1})\operatorname{\downarrow}=1$. We let $x_{0}^{*}$ be the minimum of all such $x_{0}$, and $x_{1}^{*}$ the maximum of all such $x_{1}$, across all $\sigma\in T_{e,s}$ of length $s$. (Note that there may be multiple pairs $x_{0}<x_{1}$ even for a single $\sigma$.) We then enumerate every element of $[x_{0}^{*},x_{1}^{*})$ into $S$ and say $\mathcal{P}_{e}$ has acted. Finally, we cancel the restraints of any lower priority $\mathcal{L}$ requirements. This completes the construction of $S$. We claim that each requirement acts at most finitely often. Suppose not, and fix the highest priority requirement that acts at infinitely many stages. Let $s<t$ be stages at which this requirement acts, but such that every stronger requirement has stopped acting before stage $s$. Then this requirement must require attention at stage $t$, and moreover, it must not have acted since the last stage at which some stronger $\mathcal{P}$ requirement did. In particular, the requirement must not have acted at stage $s$, a contradiction. So the claim is proved. Now, a restraint $r_{e}$ is only ever cancelled if some $\mathcal{P}$ requirement stronger than $\mathcal{L}_{e}$ acts, and it is only defined if $\mathcal{L}_{e}$ itself acts. The preceding claim thus implies that each $r_{e}$ is either undefined from some stage on or reaches a final value. From here, a straightforward induction establishes that for each $e$, $\lim_{s}T_{e,s}$ exists. Namely, $\lim_{s}T_{e,s}$ is the set of all $\sigma\in\lim_{s}T_{e-1,s}$ such that there do not exist $x_{0}<x_{1}$, larger than the final restraints imposed by any $\mathcal{L}$ requirements stronger than $\mathcal{P}_{e}$, such that $\Phi_{e}^{\sigma}(x_{0})\operatorname{\downarrow}=\Phi_{e}^{\sigma}(x_{1})\operatorname{\downarrow}=1$, provided this set is infinite, or $\lim_{s}T_{e-1,s}$ if this set is finite. Let $T_{e}=\lim_{s}T_{e-1,s}$. We have $T\supseteq T_{0}\supseteq T_{1}\supseteq\cdots$, a sequence of infinite binary trees, so by compactness of Cantor space we have that $\bigcap_{e}[T_{e}]\neq\emptyset$. Let $P_{0}$ be any element of this intersection. By choice of $T$, we have that $P_{0}\gg\emptyset^{\prime}$. It remains to verify that all requirements are satisfied. Consider any requirement, and let $s$ be least such that every higher priority $\mathcal{P}$ requirement has stopped acting prior to stage $s$. If we are dealing with $\mathcal{L}_{e}$, and if $\Phi_{e}^{L_{0}}$ is the characteristic function of an infinite set, then $\mathcal{L}_{e}$ will require attention at stage $s$. This means that for some $x_{1}>x_{0}\geq m_{s}$ with $\Phi_{e}^{L_{0}}(x_{0})\operatorname{\downarrow}=\Phi^{L_{0}}_{e}(x_{1})\operatorname{\downarrow}=1$, $\mathcal{L}_{e}$ will now impose the restraint $r_{e}=x_{1}$, and this restraint will be permanent. Only $\mathcal{P}$ requirements enumerate elements into $S$, and the only such requirements that can act after stage $s$ must respect $r_{e}$. Hence, every number in the interval $[x_{0},x_{1})$ will be permanently kept out of $S$, and so belong to $\overline{S}$, as needed. Now suppose we are dealing with $\mathcal{P}_{e}$, and that $\Phi_{e}^{P_{0}}$ is the characteristic function of an infinite set. Any $\mathcal{L}$ requirement stronger than $\mathcal{P}_{e}$ that acts at or after stage $s$ will do so before $\mathcal{P}_{e}$ can act. This is because the action of $\mathcal{P}_{e}$ cannot change any such $\mathcal{L}$ requirement from not requiring attention to requiring attention. Let $t\geq s$ be least such that all such $\mathcal{L}$ requirements have acted prior to stage $t$, so that in particular, any restraints they impose are now permanent. By hypothesis on $P_{0}$, the choice of $s$, and the definition of $T_{e}$, there is a $u\geq t$ such that the strings of length $u$ in $T_{e,u}$ are precisely the same as the strings of length $u$ in $T_{e}$, and such that $\mathcal{P}_{e}$ requires attention at stage $u$. Choose the least such $u$. Then $\mathcal{P}_{e}$ will act at stage $u$, which means that for every $\sigma\in T_{e,u}$ of length $u$ there will be some $x_{1}>x_{0}$ such that $\Phi_{e}^{\sigma}(x_{0})\operatorname{\downarrow}=\Phi^{\sigma}_{e}(x_{1})\operatorname{\downarrow}=1$ and every element of $[x_{0},x_{1})$ is enumerated into $S$. But since $P_{0}\in[T]\subseteq[T_{e}]$, it follows that $P_{0}\operatorname{\upharpoonright}u\in T_{e,u}$, so in particular there are $x_{1}>x_{0}$ such that $\Phi_{e}^{P_{0}}(x_{0})\operatorname{\downarrow}=\Phi^{P_{0}}_{e}(x_{1})\operatorname{\downarrow}=1$ and $[x_{0},x_{1})\subseteq S$. This completes the proof. ∎
@@ 7.2.1. Forcing notion for $G$
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 7.2.2. Forcing notion for $L_{G}$
Then $\mathcal{C}$ is a $\Pi^{0}_{1}(X\oplus Y)$ class, so if it is non-empty we can apply Lemma 7.8 to find a $\langle Y_{0},Y_{1}\rangle\in\mathcal{C}$ and an $i<2$ such that $(D^{0},D^{1},Y_{i})$ is a constructible extension of $(D^{0},D^{1},Y)$. Then the $L_{G}$-condition obtained from $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ by replacing $Y$ with $Y_{i}$ and $J^{i}$ with $J^{i}\cup\{e_{i}\}$ is a constructible extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ that forces $e_{i}\notin L_{\mathtt{G}}^{\prime}$ on the $i$ side. Suppose, then, that $\mathcal{C}$ is empty. By compactness of Cantor space, there is an $\ell\in\omega$ such that for every $2$-partition $\langle Y_{0},Y_{1}\rangle$ of $Y$ there exists an $i<2$ and an $F\subseteq Y_{i}\operatorname{\upharpoonright}\ell$ for which $T_{D^{i}\cup F,H^{i},J^{i}\cup\{e_{i}\}}$ does not look extendible. Fix such an $i$ and $F$ for the $2$-partition $\langle Y\cap A^{0},Y\cap A^{1}\rangle$. Then the condition obtained from $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ by replacing $D^{i}$ with $D^{i}\cup F$ and $Y$ by $Y\smallsetminus[0,\max F]$ is a finite extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ that forces $e_{i}\in L_{\mathtt{G}}^{\prime}$ on the $i$ side. ∎
@@
which is a $\Pi^{0}_{1}(X\oplus Y)$ class. If there is a $k$ for which $\mathcal{C}_{k}$ is non-empty we apply Lemma 7.8 to find a $\langle Y_{0},Y_{1}\rangle\in\mathcal{C}_{k}$ and an $i<2$ such that $(D^{0},D^{1},Y_{i})$ is a constructible extension of $(D^{0},D^{1},Y)$. The $L_{G}$-condition obtained by replacing $Y$ with $Y_{i}$ and $H^{i}$ with $H^{i}\cup\{\langle e_{i},k\rangle\}$ is then a constructible extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ that forces $\Phi_{e_{i}}^{L_{\mathtt{G}}}$ is finite on the $i$ side. Suppose next that $\mathcal{C}_{k}$ is empty for every $k$. This means that for every $k$, there is an $\ell$ such that for every $2$-partition $\langle Y_{0},Y_{1}\rangle$ of $Y$ there is an $i<2$ and a finite set $F\subseteq Y_{i}\operatorname{\upharpoonright}\ell$ such that $T_{D^{i}\cup F,H^{i}\cup\{\langle e_{i},k\rangle\},J^{i}}$ does not look extendible. Let $\ell_{k}$ be the least such $\ell$. Since $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ is a condition and $F\subseteq Y_{i}\subseteq Y$, the tree $T_{D^{1}\cup F,H^{1},J^{1}}$ looks extendible, and so contains a string $\sigma$ of maximal length. Since no such $\sigma$ belongs to $T_{D^{i}\cup F,H^{i}\cup\{\langle e_{i},k\rangle\},J^{i}}$, it follows from the definition that there exist $x_{0},x_{1}\leq|\sigma|$ such that $x_{1}>x_{0}>k$ and $\Phi_{e_{i}}^{\sigma}(x_{0})\operatorname{\downarrow}=\Phi_{e_{1}}^{\sigma}(x_{1})\operatorname{\downarrow}=1$. Let $x^{k}_{0}$ be the minimum of all such $x_{0}$, and $x^{k}_{1}$ the maximum of all the such $x_{1}$, across both $i<2$, all finite sets $F\subseteq Y_{i}\operatorname{\upharpoonright}\ell_{k}$ as above, and all $\sigma\in T_{D^{i}\cup F,H^{i},J^{i}}$ of maximal length. Now, $\ell_{k}$ can be found uniformly $(X\oplus Y)$-computably from $k$, and hence so can the finite collection of all the finite sets $F$ above. It follows that $X\oplus Y$ can uniformly compute the numbers $x^{k}_{0}$ and $x^{k}_{1}$ for each $k$. Since $S$ is $L$-hyperimmune, and since $L\geq_{\text{T}}X\oplus Y$ and $k<x^{k}_{0}<x^{k}_{1}$ for all $k$, there must be a $k$ such that $[x^{k}_{0},x^{k}_{1})\subseteq\overline{S}$. Fix this $k$, and fix $i<2$ and $F\subseteq(Y\cap A^{i})\operatorname{\upharpoonright}\ell_{k}$ such that $T_{D^{i}\cup F,H^{i}\cup\{\langle e_{i},k\rangle\},J^{i}}$ does not look extendible. Then the condition obtained from $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ by replacing $D^{i}$ by $D^{i}\cup F$ and $Y$ by $Y\smallsetminus[0,\max F]$ is a finite extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ that forces $\Phi_{e_{i}}^{L_{\mathtt{G}}}$ is not a non-hyperimmunity witness for $S$ on the $i$ side. ∎
@@
which is a $\Pi^{0}_{1}(X\oplus Y)$ class. If there is a $k$ for which $\mathcal{C}_{k}$ is non-empty we apply Lemma 7.8 to find a $\langle Y_{0},Y_{1}\rangle\in\mathcal{C}_{k}$ and an $i<2$ such that $(D^{0},D^{1},Y_{i})$ is a constructible extension of $(D^{0},D^{1},Y)$. The $L_{G}$-condition obtained by replacing $Y$ with $Y_{i}$ and either $J^{0}$ with $J^{0}\cup\{e_{0}\}$ if $i=0$, or $H^{1}$ with $H^{1}\cup\{\langle e_{1},k\rangle\}$ if $i=1$, is then a constructible extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$. If $i=0$ this extension forces $e_{0}\notin L_{\mathtt{G}}^{\prime}$ on the 00 side, and if $i=1$ it forces $\Phi_{e_{1}}^{L_{\mathtt{G}}}$ is finite on the $1$ side. Now suppose $\mathcal{C}_{k}$ is empty for every $k$. This means there is an $\ell_{k}$ such that for every $2$-partition $\langle Y_{0},Y_{1}\rangle$ of $Y$, one of the following holds:
@@
    A $P_{G}$-precondition is a tuple $(p,b^{0},b^{1})$ such that $p$ is an $L_{G}$ condition, $b^{0},b^{1}\in\omega$, and for each $i<2$ and each $e<b^{i}$, $p$ either forces $e\in L_{\mathtt{G}}^{\prime}$ or $e\notin L_{\mathtt{G}}^{\prime}$ on the $i$ side.
@@
    A precondition $(q,c^{0},c^{1})$ extends $(p,b^{0},b^{1})$ if $q$ extends $p$ as $L_{G}$-conditions, and if for each $i<2$, $c^{i}\geq b^{i}$. $q$ is a constructible extension if $q$ is a constructible extension of $p$ as $L_{G}$-conditions.
@@
    A condition $(q,c^{0},c^{1},W^{0},W^{1})$ extends $(p,b^{0},b^{1},V^{0},V^{1})$ if $q$ extends $p$ as $L_{G}$-conditions, and if for each $i<2$, $c^{i}\geq b^{i}$ and $V^{i}\subseteq W^{i}$. $q$ is a constructible extension if $q$ is a constructible extension of $p$ as $L_{G}$-conditions.

### v2

@@ 7. Separating $\mathsf{GST}_{1}$ from $\mathsf{RT}^{2}_{2}$
We call requirements of the former kind the $\mathcal{R}$ requirements, and those of the latter kind the $\mathcal{Q}$ requirements. It is clear that satisfying all $\mathcal{R}$ and $\mathcal{Q}$ requirements ensures that $S$ will be $L_{0}$-hyperimmune and $\overline{S}$ will be $L_{0}^{\prime}$-hyperimmune.
We assign a priority ordering to the requirements of type $\omega$, with $\mathcal{Q}_{e}$ stronger than $\mathcal{R}_{e}$, and $\mathcal{R}_{e}$ stronger than $\mathcal{Q}_{e+1}$, for all $e\in\omega$. Requirements will act at certain stages, never more than one requirement per stage, in a way that we will make precise below. When some $\mathcal{R}_{e}$ acts, it will impose a restraint $r_{e}$ that may be cancelled later and redefined later still; this will be used to control which elements can be enumerated into $S$ by lower priority $\mathcal{Q}$ requirements. At each stage $s$, we also let $m_{s}$ be the maximum of all restraints (including cancelled or redefined ones) imposed at any point in the construction prior to stage $s$, as well as all numbers enumerated into $S$ so far.
We say a requirement $\mathcal{R}_{e}$ requires attention at stage $s$ if there exist numbers $x_{1}>x_{0}\geq m_{s}$ such that $\Phi_{e}^{L_{0}}(x_{0})\operatorname{\downarrow}=\Phi_{e}^{L_{0}}(x_{1})\operatorname{\downarrow}=1$. We say a requirement $\mathcal{Q}_{e}$ requires attention at stage $s$ if there exist numbers $x_{0}<x_{1}\leq s$, larger than all active restraints at stage $s$ imposed by any $\mathcal{R}$ requirements stronger than $\mathcal{Q}_{e}$, such that $\Phi_{e}^{L_{0}^{\prime}}(x_{0})[s]\operatorname{\downarrow}=\Phi_{e}^{L_{0}^{\prime}}(x_{1})[s]\operatorname{\downarrow}=1$. Notice that, in either case, determining if a given requirement requires attention at a given stage is computable in $L^{\prime}_{0}$, uniformly in the index of the requirement and the stage number.
Initially, we let $S$ be empty and let no restraints be imposed. (In particular, $m_{0}=0$.) Now suppose we are at stage $s$. If it exists, fix the highest priority requirement of index $e\leq s$ that requires attention at stage $s$ and that has not acted since the last stage, if any, at which a higher priority $\mathcal{Q}$ requirement acted. If this requirement is $\mathcal{R}_{e}$, we fix the least witnessing $x_{1}>x_{0}\geq m_{s}$ and impose the restraint $r_{e}=x_{1}$. We then say that $\mathcal{L}_{e}$ has acted. If, instead, the requirement is $\mathcal{Q}_{e}$, we fix the least witnessing $x_{0}<x_{1}$ and enumerate every element of the interval $[x_{0},x_{1})$ into $S$. We then say $\mathcal{Q}_{e}$ has acted, and cancel the restraints of any lower priority $\mathcal{R}$ requirements.
This completes the construction. We claim that each requirement acts at most finitely often. Suppose not, and fix the highest priority requirement that acts at infinitely many stages. Let $s<t$ be stages at which this requirement acts, but such that every stronger requirement has stopped acting before stage $s$. Then this requirement must require attention at stage $t$, and moreover, it must not have acted since the last stage at which some stronger $\mathcal{Q}$ requirement did. In particular, the requirement must not have acted at stage $s$, a contradiction. So the claim is proved.
Now, a restraint $r_{e}$ is only ever cancelled if some $\mathcal{Q}$ requirement stronger than $\mathcal{R}_{e}$ acts, and it is only defined if $\mathcal{R}_{e}$ itself acts. The preceding claim thus implies that each $r_{e}$ is either undefined from some stage on or reaches a final value.
It remains to verify that all requirements are satisfied. Consider any requirement, and let $s$ be least such that every higher priority $\mathcal{Q}$ requirement has stopped acting prior to stage $s$. If we are dealing with $\mathcal{R}_{e}$, and if $\Phi_{e}^{L_{0}}$ is the characteristic function of an infinite set, then $\mathcal{R}_{e}$ will require attention at stage $s$. This means that for some $x_{1}>x_{0}\geq m_{s}$ with $\Phi_{e}^{L_{0}}(x_{0})\operatorname{\downarrow}=\Phi^{L_{0}}_{e}(x_{1})\operatorname{\downarrow}=1$, $\mathcal{R}_{e}$ will now impose the restraint $r_{e}=x_{1}$, and this restraint will be permanent. Only $\mathcal{Q}$ requirements enumerate elements into $S$, and the only such requirements that can act after stage $s$ must respect $r_{e}$. Hence, every number in the interval $[x_{0},x_{1})$ will be permanently kept out of $S$, and so belong to $\overline{S}$, as needed.
Now suppose we are dealing with $\mathcal{Q}_{e}$, and that $\Phi_{e}^{L_{0}^{\prime}}$ is the characteristic function of an infinite set. Any $\mathcal{R}$ requirement stronger than $\mathcal{Q}_{e}$ that acts at or after stage $s$ will do so before $\mathcal{Q}_{e}$ can act. This is because the action of $\mathcal{Q}_{e}$ cannot change any such $\mathcal{R}$ requirement from not requiring attention to requiring attention. Let $t\geq s$ be least such that all such $\mathcal{R}$ requirements have acted prior to stage $t$. In particular, any restraints they impose are now permanent, by choice of $s$. By hypothesis, there is a $u\geq t$ such that $\mathcal{Q}_{e}$ acts at stage $u$, which means that for some $x_{1}>x_{0}$ such that $\Phi_{e}^{L_{0}^{\prime}}(x_{0})\operatorname{\downarrow}=\Phi^{L_{0}^{\prime}}_{e}(x_{1})\operatorname{\downarrow}=1$ every element of $[x_{0},x_{1})$ is enumerated into $S$. This completes the proof. ∎
@@ 7.2.1. Forcing notion for $G$
    An index for $(D^{0},D^{1},Y)$ is a number $\langle d_{0},d_{1},y\rangle$ such that for each $i<2$, $d_{i}$ is a canonical index for $D^{i}$, and $\Phi^{L}_{y}$ is a DNC2 function relative to $X\oplus Y$.
@@
From an index for $(D^{0},D^{1},Y)$, we first extract a $y\in\omega$ such that $\Phi^{L}_{y}$ is DNC2 relative to $X\oplus Y$. From a canonical index for $Y\smallsetminus Z$, we can uniformly find a computable function $h$ such that $\Phi^{X\oplus Y}_{h(e)}(x)=\Phi^{X\oplus Z}_{e}(e)$ for all $e$ and $x$. Now let $g$ be a computable function so that $\Phi^{L}_{g(e)}(x)=\Phi^{L}_{e}(h(x))$ for all $e$ and $x$. Then for all $e$ we have that
@@
so $\Phi^{L}_{g(y)}$ is DNC2 relative to $X\oplus Z$. Thus, the canonical indices of $E^{0}$ and $E^{1}$, together with $g(y)$, form an index for $(E^{0},E^{1},Z)$. ∎
@@
First, from an index for $(D^{0},D^{1},Y)$, fix $y\in\omega$ such that $\Phi^{L}_{y}$ is DNC2 relative to $X\oplus Y$. Let $\mathcal{D}$ be the class of all $\langle f_{0},f_{1},Y_{0},Y_{1}\rangle$ such that $\langle Y_{0},Y_{1}\rangle\in\mathcal{C}$ and, for each $i<2$, $f_{i}$ is DNC2 over $X\oplus Y_{i}$. This is a non-empty $\Pi^{0}_{1}(X\oplus Y)$ class, whose index as such can be found uniformly computably from such an index for $\mathcal{C}$. Thus, $\Phi^{L}_{y}$ can uniformly compute an element $\langle f_{0},f_{1},Y_{0},Y_{1}\rangle\in\mathcal{D}$. In particular, this produces indices $y_{i}$ for each $i<2$ such that $\Phi_{y_{i}}^{L}=f_{i}$. Now, since $Y$ is infinite, there is at least one $i<2$ such that $Y_{i}$ is infinite, and since $P\gg L^{\prime}$ it follows that $P$ can uniformly compute such an $i$ (see [16], Theorem 2.8.25 (4)). Thus, the canonical indices for $D^{0}$ and $D^{1}$, together with $y_{i}$, form an index for $(D^{0},D^{1},Y_{i})$. ∎
@@ 7.2.2. Forcing notion for $L_{G}$
Then $\mathcal{C}$ is a $\Pi^{0}_{1}(X\oplus Y)$ class, so if it is non-empty we can apply Lemma 7.8 to find a $\langle Y_{0},Y_{1}\rangle\in\mathcal{C}$ and an $i<2$ such that $(D^{0},D^{1},Y_{i})$ is a constructible extension of $(D^{0},D^{1},Y)$. Then the $L_{G}$-condition obtained from $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ by replacing $Y$ with $Y_{i}$ and $J^{i}$ with $J^{i}\cup\{e_{i}\}$ is a constructible extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ that forces $e_{i}\notin L_{\mathtt{G}}^{\prime}$ on the $i$ side.
Suppose, then, that $\mathcal{C}$ is empty. By compactness of Cantor space, there is an $\ell\in\omega$ such that for every $2$-partition $\langle Y_{0},Y_{1}\rangle$ of $Y$ there exists an $i<2$ and an $F\subseteq Y_{i}\operatorname{\upharpoonright}\ell$ for which $T_{D^{i}\cup F,H^{i},J^{i}\cup\{e_{i}\}}$ does not look extendible. Fix such an $i$ and $F$ for the $2$-partition $\langle Y\cap A^{0},Y\cap A^{1}\rangle$. Then the condition obtained from $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ by replacing $D^{i}$ with $D^{i}\cup F$ and $Y$ by $Y\smallsetminus[0,\max F]$ is a finite extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ that forces $e_{i}\in L_{\mathtt{G}}^{\prime}$ on the $i$ side. ∎
@@
which is a $\Pi^{0}_{1}(X\oplus Y)$ class. If there is a $k$ for which $\mathcal{C}_{k}$ is non-empty we apply Lemma 7.8 to find a $\langle Y_{0},Y_{1}\rangle\in\mathcal{C}_{k}$ and an $i<2$ such that $(D^{0},D^{1},Y_{i})$ is a constructible extension of $(D^{0},D^{1},Y)$. The $L_{G}$-condition obtained by replacing $Y$ with $Y_{i}$ and $H^{i}$ with $H^{i}\cup\{\langle e_{i},k\rangle\}$ is then a constructible extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ that forces $\Phi_{e_{i}}^{L_{\mathtt{G}}}$ is finite on the $i$ side.
Suppose next that $\mathcal{C}_{k}$ is empty for every $k$. This means that for every $k$, there is an $\ell$ such that for every $2$-partition $\langle Y_{0},Y_{1}\rangle$ of $Y$ there is an $i<2$ and a finite set $F\subseteq Y_{i}\operatorname{\upharpoonright}\ell$ such that $T_{D^{i}\cup F,H^{i}\cup\{\langle e_{i},k\rangle\},J^{i}}$ does not look extendible. Let $\ell_{k}$ be the least such $\ell$. Since $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ is a condition and $F\subseteq Y_{i}\subseteq Y$, the tree $T_{D^{1}\cup F,H^{1},J^{1}}$ looks extendible, and so contains a string $\sigma$ of maximal length. Since no such $\sigma$ belongs to $T_{D^{i}\cup F,H^{i}\cup\{\langle e_{i},k\rangle\},J^{i}}$, it follows from the definition that there exist $x_{0},x_{1}\leq|\sigma|$ such that $x_{1}>x_{0}>k$ and $\Phi_{e_{i}}^{\sigma}(x_{0})\operatorname{\downarrow}=\Phi_{e_{1}}^{\sigma}(x_{1})\operatorname{\downarrow}=1$. Let $x^{k}_{0}$ be the minimum of all such $x_{0}$, and $x^{k}_{1}$ the maximum of all the such $x_{1}$, across both $i<2$, all finite sets $F\subseteq Y_{i}\operatorname{\upharpoonright}\ell_{k}$ as above, and all $\sigma\in T_{D^{i}\cup F,H^{i},J^{i}}$ of maximal length.
Now, $\ell_{k}$ can be found uniformly $(X\oplus Y)$-computably from $k$, and hence so can the finite collection of all the finite sets $F$ above. It follows that $X\oplus Y$ can uniformly compute the numbers $x^{k}_{0}$ and $x^{k}_{1}$ for each $k$. Since $S$ is $L$-hyperimmune, and since $L\geq_{\text{T}}X\oplus Y$ and $k<x^{k}_{0}<x^{k}_{1}$ for all $k$, there must be a $k$ such that $[x^{k}_{0},x^{k}_{1})\subseteq\overline{S}$. Fix this $k$, and fix $i<2$ and $F\subseteq(Y\cap A^{i})\operatorname{\upharpoonright}\ell_{k}$ such that $T_{D^{i}\cup F,H^{i}\cup\{\langle e_{i},k\rangle\},J^{i}}$ does not look extendible. Then the condition obtained from $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ by replacing $D^{i}$ by $D^{i}\cup F$ and $Y$ by $Y\smallsetminus[0,\max F]$ is a finite extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$ that forces $\Phi_{e_{i}}^{L_{\mathtt{G}}}$ is not a non-hyperimmunity witness for $S$ on the $i$ side. ∎
@@
which is a $\Pi^{0}_{1}(X\oplus Y)$ class. If there is a $k$ for which $\mathcal{C}_{k}$ is non-empty we apply Lemma 7.8 to find a $\langle Y_{0},Y_{1}\rangle\in\mathcal{C}_{k}$ and an $i<2$ such that $(D^{0},D^{1},Y_{i})$ is a constructible extension of $(D^{0},D^{1},Y)$. The $L_{G}$-condition obtained by replacing $Y$ with $Y_{i}$ and either $J^{0}$ with $J^{0}\cup\{e_{0}\}$ if $i=0$, or $H^{1}$ with $H^{1}\cup\{\langle e_{1},k\rangle\}$ if $i=1$, is then a constructible extension of $(D^{0},D^{1},Y,H^{0},J^{0},H^{1},J^{1})$. If $i=0$ this extension forces $e_{0}\notin L_{\mathtt{G}}^{\prime}$ on the 00 side, and if $i=1$ it forces $\Phi_{e_{1}}^{L_{\mathtt{G}}}$ is finite on the $1$ side.
Now suppose $\mathcal{C}_{k}$ is empty for every $k$. This means there is an $\ell_{k}$ such that for every $2$-partition $\langle Y_{0},Y_{1}\rangle$ of $Y$, one of the following holds:
@@
    A $P_{G}$-precondition is a tuple $(p,b^{0},b^{1})$ such that $p$ is an $L_{G}$-condition, $b^{0},b^{1}\in\omega$, and for each $i<2$ and each $e<b^{i}$, $p$ either forces $e\in L_{\mathtt{G}}^{\prime}$ or $e\notin L_{\mathtt{G}}^{\prime}$ on the $i$ side.
@@
    A precondition $(q,c^{0},c^{1})$ extends $(p,b^{0},b^{1})$ if $q$ extends $p$ as $L_{G}$-conditions, and if for each $i<2$, $c^{i}\geq b^{i}$. It is a constructible extension if $q$ is a constructible extension of $p$ as $L_{G}$-conditions.
@@
    A condition $(q,c^{0},c^{1},W^{0},W^{1})$ extends $(p,b^{0},b^{1},V^{0},V^{1})$ if $q$ extends $p$ as $L_{G}$-conditions, and if for each $i<2$, $c^{i}\geq b^{i}$ and $V^{i}\subseteq W^{i}$. It is a constructible extension if $q$ is a constructible extension of $p$ as $L_{G}$-conditions.